### v1

# LLM-guided Chemical Process Optimization with a Multi-Agent Approach

Abstract
Chemical process optimization is crucial to maximize production efficiency and economic performance. Traditional methods, including gradient-based solvers, evolutionary algorithms, and parameter grid searches, become impractical when operating constraints are ill-defined or unavailable, requiring engineers to rely on subjective heuristics to estimate feasible parameter ranges. To address this constraint definition bottleneck, we present a multi-agent framework of large language model (LLM) agents that autonomously infer operating constraints from minimal process descriptions, then collaboratively guide optimization using the inferred constraints. Our AutoGen-based agentic framework employs OpenAI’s o3 model, with specialized agents for constraint generation, parameter validation, simulation execution, and optimization guidance. Through two phases - autonomous constraint generation using embedded domain knowledge, followed by iterative multi-agent optimization - the framework eliminates the need for predefined operational bounds. Validated on the hydrodealkylation process across cost, yield, and yield-to-cost ratio metrics, the framework demonstrated competitive performance with conventional optimization methods while achieving better computational efficiency, requiring fewer iterations to converge. Our approach converged in under 20 minutes, achieving a 31-fold speedup over grid search. Beyond computational efficiency, the framework’s reasoning-guided search demonstrates sophisticated process understanding, correctly identifying utility trade-offs, and applying domain-informed heuristics. This approach shows significant potential for optimization scenarios where operational constraints are poorly characterized or unavailable, particularly for emerging processes and retrofit applications.
[…]
{tocentry}[Uncaptioned image]
1 Introduction
Efficient optimization of chemical processes is a central objective in process systems engineering, with direct implications for production efficiency and economic performance1. Conventional optimization approaches include gradient-based solvers such as Interior Point OPTimizer (IPOPT)2 and Sparse Nonlinear OPTimizer (SNOPT)3, which depend on the smoothness and differentiability of the underlying objective function4. When these requirements cannot be met, alternative methods such as evolutionary algorithms5, and derivative-free methods6 provide viable solutions. Furthermore, exhaustive search techniques7 such as grid search systematically evaluate all combinations of discretized parameter values throughout the design space.
However, these optimization methods require a complete and well-defined design space with described operating constraints. This requirement makes it difficult to apply these methods when feasible parameter ranges or process operating conditions are incomplete or unavailable. In these cases, engineers typically rely on heuristics or domain knowledge to estimate feasible ranges8, 9. These heuristic approaches are highly subjective and may lead to variable optimization results. Mathematical simulation platforms such as IDAES10 and modeling libraries such as Pyomo11 offer flexible environments to build detailed process models, but still depend on users to define variable constraints and smooth objective functions before running optimization. This constraint definition bottleneck significantly limits the application of such optimization techniques, particularly for non-expert users or for novel processes where operational configurations are not well-established.
Recent developments in large language models (LLMs) offer potential approaches to address these limitations. LLMs have demonstrated strong reasoning capabilities in complex problem-solving tasks12, 13 and effective adaptation to previously unseen domains through few-shot learning14. Most remarkably, recent work has shown that LLMs can autonomously conduct scientific research, from hypothesis generation and experimental design to data analysis and manuscript writing15, demonstrating their potential for independent scientific reasoning and discovery.
These general capabilities have translated into significant advances within chemical engineering applications. In synthesis planning, LLMs evaluate synthetic routes based on strategic requirements16, suggesting that domain expertise traditionally requiring years of training could become more accessible through natural language interfaces. For material discovery, LLM-based agents such as ChemCrow17 have successfully planned and executed syntheses using automated laboratory platforms. For control system design18, 19, 20, 21, 22, LLMs enable more flexible process control approaches without the need for specialized programming skills.
However, most existing efforts focus on isolated decision steps or downstream control, with limited exploration of LLM integration into steady-state simulation or process optimization pipelines, and autonomous constraint generation for steady-state process optimization using collaborative multi-agent LLM systems remains largely unexplored. This work addresses this gap by introducing an autonomous, multi-agent framework that integrates language model reasoning with process simulation for operating condition optimization. We tackle the challenge of autonomous constraint generation when feasible operating ranges are unknown or poorly defined.
Our approach leverages the AutoGen platform23 with OpenAI’s o3 model24 to create a multi-agent system where language model agents collaborate to both generate realistic constraints and iteratively optimize process parameters. The framework is evaluated on a real-world chemical flowsheet - the hydrodealkylation (HDA) process using IDAES10 simulations and assessed across three metrics: cost, yield, and yield-to-cost ratio. We selected IPOPT and grid search as representative benchmarks: IPOPT as a state-of-the-art gradient-based solver widely used in chemical engineering, and grid search as a gradient-free exhaustive method that serves as a baseline for global optimization performance. The results show comparable performance, while requiring fewer iterations and less computational time. This approach could significantly reduce the expertise barrier for process optimization and enable more widespread application of advanced optimization techniques in industrial settings, particularly for emerging processes or retrofitting existing operations where operational constraints are poorly characterized.
2 Methodology
2.1 Framework Architecture
We propose a multi-agent interactive system using AutoGen’s GroupChat architecture that addresses fundamental limitations of single-agent approaches for complex optimization tasks. Individual agents maintain focused prompts for specific functions, preventing context dilution that occurs when lengthy multi-purpose prompts overwhelm LLM attention mechanisms. Each agent’s specialized role enables direct function calling rather than complex output parsing, reducing implementation complexity and error potential. The modular design allows independent debugging and modification of optimization components, critical for iterative algorithm development and deployment flexibility.
Our framework consists of five specialized LLM agents, where each "agent" represents a specialized LLM instance with distinct prompts and function-calling capabilities designed for specific optimization tasks. The system operates through two phases: constraint generation followed by iterative optimization within a collaborative multi-agent environment, as shown in Figure 1. Throughout this process, autonomy refers to each agent’s ability to reason, act, and iterate without human intervention, using only embedded domain knowledge, generated constraints, and shared optimization history.
Figure 1: Architecture of the LLM-based optimization framework illustrating agent workflow. The ContextAgent generates constraints and detailed process overview from basic process info using domain knowledge, while specialized agents within the GroupChat environment handle parameter introduction, validation, optimization suggestions, and IDAES-based simulation evaluation through iterative refinement until optimal solutions are achieved.
The first phase involves the ContextAgent operating independently to infer realistic variable bounds and generate process context from given descriptions using embedded domain knowledge. This agent infers constraint ranges as lower and upper bounds for each decision variable and creates a detailed process overview text that provides contextual understanding for subsequent optimization agents (detailed process overview generated by ContextAgent shown in Supporting Information). These generated constraints and process overview serve as the foundation for the subsequent optimization phase.
The second phase uses the GroupChat environment to support coordinated interactions and shared memory between four specialized agents during optimization. Through iterative cycles of parameter proposal, validation, simulation, and refinement, the agents collaboratively navigate the constraint-defined parameter space toward optimal solutions.
2.2 Agent Roles and Functions
Within the GroupChat architecture, each agent maintains specialized responsibilities that coordinate to enable autonomous optimization. The optimization process operates through coordinated agent interactions in a specific sequence:
ParameterAgent initiates the process by receiving parameter-value pairs from users as initial starting points. These initial guesses can be arbitrary and may violate feasible bounds, as users typically lack knowledge of the constraint ranges previously generated by the ContextAgent. The ValidationAgent will identify any constraint violations and redirect to the SuggestionAgent for correction.
ValidationAgent serves as the critical first checkpoint, evaluating each parameter set against the generated constraints and providing feedback about whether proposed values are acceptable or identifying which specific parameters need adjustment due to constraint violations. When invalid input is detected, the system automatically redirects to the SuggestionAgent for revision, creating a validation-correction loop that ensures all parameter sets remain within feasible bounds.
SimulationAgent executes process evaluation once parameters pass validation. This agent runs a custom objective function that interfaces with a pre-defined python-based IDAES simulation model (detailed IDAES implementation, thermodynamic properties, and kinetic parameters are provided in Supporting Information), calculating performance metrics such as cost, yield, and yield-to-cost ratios for each parameter configuration. Performance results are reported back to the SuggestionAgent, including any simulation convergence issues.
SuggestionAgent serves as the optimization engine, maintaining a comprehensive record of the optimization history. After each iteration, current parameter values and corresponding simulation results are appended to the agent’s memory, providing direct access to both valid bounds and historical trial data. This accumulated knowledge enables the agent to refine and propose future parameter sets based on observed trends and constraint violations, completing the iterative optimization cycle.
Agent behavior is controlled through carefully designed prompts that define roles, objectives, and output formats. The ContextAgent prompt incorporates engineering domain knowledge and formatting requirements to generate realistic operating constraints from process descriptions, while the SuggestionAgent prompt implements memory-guided optimization logic with constraint violation handling. Complete prompt templates are provided in Supporting Information, demonstrating the integration of process engineering heuristics with structured output requirements that enable autonomous constraint generation and iterative parameter refinement.
2.3 Optimization Workflow
The iterative optimization process operates through a precisely defined agent interaction sequence that ensures systematic parameter refinement. Agent interactions are coordinated through a customized selector function (detailed algorithm is provided in Supporting Information) that determines speaking order based on current speaker and message content, preventing redundant communications and ensuring logical workflow progression.
The core optimization cycle follows this sequence: ParameterAgent introduces values $\rightarrow$ ValidationAgent checks feasibility $\rightarrow$ SimulationAgent evaluates performance $\rightarrow$ SuggestionAgent analyzes results and proposes improvements $\rightarrow$ ValidationAgent verifies new proposals. This cycle repeats autonomously until the SuggestionAgent detects convergence based on diminishing performance improvements across successive iterations.
This structured workflow design addresses three critical challenges in autonomous optimization: (1) constraint enforcement through mandatory validation checkpoints, (2) systematic exploration through memory-guided suggestions, and (3) autonomous termination through trend analysis. Unlike gradient-based methods that require mathematical derivatives or grid search that evaluates fixed discretizations, this agent-based approach leverages reasoning to intelligently navigate parameter spaces, adapting search strategy based on constraint violations and performance feedback. The framework’s efficiency stems from this reasoning-guided approach, enabling convergence in 20-45 iterations compared to thousands required by exhaustive methods, while maintaining solution quality through systematic constraint enforcement and trend analysis.
The framework’s autonomous nature is demonstrated through self-determined termination when the SuggestionAgent detects diminishing returns based on simulation history. This approach reflects the agent’s ability to prevent unnecessary computation while adapting to different problem characteristics, ensuring computational efficiency without sacrificing solution quality.
2.4 Evaluation Metrics
The SimulationAgent evaluates process performance using three metrics that capture the primary objectives of chemical process optimization: operating cost, product yield, and yield-to-cost ratio.
Annual operating cost serves as the primary economic indicator, minimizing energy consumption through optimal utility management including heating costs and cooling costs for operating units:
[…]
where the detailed heating and cooling cost calculations with utility coefficients are provided in Supporting Information.
Annual product yield represents production efficiency, maximizing production capacity by optimizing the absolute production rate:
[…]
The yield-to-cost ratio provides a composite metric that balances these competing objectives, enabling evaluation of trade-offs between maximizing output and minimizing expenses:
[…]
These metrics were selected to cover industrial optimization priorities: economic performance (cost), production effectiveness (yield), and overall process efficiency (yield-to-cost ratio), allowing direct comparison with conventional optimization approaches across multiple performance dimensions.
3 Results and Discussion
3.1 Case Study Validation
To validate our multi-agent framework’s constraint generation and optimization capabilities, we demonstrate its performance using the HDA process, which shows the constraint-limited optimization challenge our approach addresses. This process converts toluene to benzene through connected unit operations, but as shown in Figure 2, only basic information is available: feed conditions, conversion rates, and equipment configuration. No operating temperature ranges, pressure limits, or feasible parameter bounds are specified. This process represents a realistic industrial scenario where detailed operational constraints are often unavailable, making it an ideal test case for autonomous constraint generation. The optimization challenge becomes apparent when attempting to apply traditional methods to this process.
Figure 2: Flowsheet of the hydrodealkylation (HDA) process used for IDAES simulation and optimization study. The process consists of mixer M101, heater H101, reactor R101, flash separators F101 and F102, splitter S101, and compressor C101. Feed streams include hydrogen (0.30 mol/s), methane (0.02 mol/s) and toluene (0.30 mol/s) at 303.15 K and 350,000 Pa. The reactor operates at 350,000 Pa with 75% toluene conversion, followed by separation units that produce benzene and toluene product streams, with 20% of the overhead stream purged through splitter S101.
Traditional optimization methods like IPOPT or grid search cannot be directly applied without explicit constraint sets. This limitation demonstrates the practical need for our ContextAgent, which infers variable bounds using embedded process knowledge. Given this constraint generation capability, we formulate the optimization problem as follows:
$\text{optimize }\phi(x)\quad\text{subject to}\quad x_{i}^{L}\leq x_{i}\leq x_{i}^{U},\quad\forall i=1,\dots,4$ (4)
where $\phi(x)$ denotes the objective function (e.g. yield, cost, or yield-to-cost ratio) evaluated by IDAES simulations, while $x^{L}$ and $x^{U}$ are variable bounds derived from LLM-generated domain knowledge. The vector $x=(x_{1},x_{2},x_{3},x_{4})$ represents the four key decision variables: H101 outlet temperature, F101 outlet temperature, F102 outlet temperature, and F102 pressure drop. The implementation of the IDAES model uses ideal vapor-liquid equilibrium relationships for thermodynamic calculations. The reactor operates with fixed 75% toluene conversion using stoichiometric reaction constraints (complete process description, feed specifications, and definitions of objective functions are provided in Supporting Information).
3.2 Constraint Generation Quality
To assess the reliability and consistency of autonomous constraint generation, the ContextAgent’s constraint generation capability was evaluated across five independent trials using identical prompts to ensure representative constraint generation and minimize stochastic variance. Table 1 presents the individual trial results and average constraint ranges inferred for the operating conditions across key unit operations H101, F101, and F102. The average ranges shown in the rightmost column were used for subsequent optimization performance comparisons.
Table 1: ContextAgent Constraint Generation Results and Average Ranges Across Five Trials
Unit Variable Trial 1 Trial 2 Trial 3 Trial 4 Trial 5 Average Range
H101 Temp (K) [870, 950] [873, 973] [820, 980] [800, 950] [773, 1023] [827.2, 975.2]
F101 Temp (K) [310, 330] [300, 340] [290, 390] [330, 420] [298, 368] [305.6, 369.6]
F102 Temp (K) [300, 340] [330, 380] [280, 340] [300, 360] [323, 413] [306.6, 366.6]
F102 $\Delta P$ (Pa) [-200000, -50000] [-200000, -50000] [-240000, -10000] [-200000, -10000] [-220000, -20000] [-216000, -28000]
As expected from the stochastic nature of LLM responses, constraint generation showed variability across trials, as detailed in the Supporting Information. For the critical H101 reactor temperature, bounds ranged from [773, 1023] K to [870, 950] K across trials, with range widths varying from 80 K to 250 K. Flash separator temperatures showed similar variability, with F101 bounds spanning from [290, 390] K to [330, 420] K, and F102 bounds from [280, 340] K to [323, 413] K. Pressure drop constraints were more consistent, with all trials generating similar lower bounds (-240000 to -200000 Pa) but varying upper bounds (-10000 to -50000 Pa). While this variability might initially appear concerning, it actually demonstrates the framework’s robust constraint generation capability. The key insight is not perfect consistency, but rather the consistent generation of industrially viable bounds that enable optimization to proceed when traditional methods cannot be applied due to missing constraint information.
Importantly, despite this observable variation across trials, the ContextAgent consistently generated realistic and engineering-appropriate bounds. For the critical H101 reactor temperature, the average generated range of [827.2, 975.2] K shows a good alignment with industrial practices reported at [773, 977] K25, 8. Although specific industrial ranges for the flash separator temperatures and pressure drop are not readily available in the literature, the close correspondence between agent-generated reactor constraints and established industrial practice demonstrates the agent’s ability to extract field-consistent operational conditions using only basic process information and embedded domain knowledge.
3.3 Performance Comparison
Optimization Results: To assess the effectiveness of our approach, we compared the LLM framework’s optimization performance against IPOPT (detailed algorithm description is provided in Supporting Information) and grid search benchmarks using the averaged ContextAgent-generated constraints. Grid search was implemented with 10 discretization points per variable, resulting in 10,000 total parameter combinations (10×10×10×10) evaluated across the four decision variables. Table 2 summarizes the optimization results across the three evaluation metrics as defined earlier.
Table 2: Comparison of IPOPT, LLM and Grid Search Optimization Results.
Metric Method H101 Temp (K) F101 Temp (K) F102 Temp (K) $\Delta P$ F102 (Pa) Value
Cost (\$/yr)a LLM 830.0 325.0 360.0 -180000 $5.792\times 10^{5}$
IPOPT 827.2 305.6 306.6 -28314 $5.660\times 10^{5}$
Grid Search 827.2 305.6 306.6 -48444 $5.660\times 10^{5}$
Yield (mol/yr)b LLM 975.0 369.5 366.0 -180000 $1.644\times 10^{7}$
IPOPT 857.5 369.6 344.4 -130340 $1.565\times 10^{7}$
Grid Search 958.8 369.6 339.9 -89333 $1.750\times 10^{7}$
Yield/Cost (mol/\$)b LLM 827.2 369.0 306.7 -28000 17.23
IPOPT 827.2 369.6 306.6 -121039 17.62
Grid Search 827.2 369.6 306.6 -68889 17.62
a Lower is better. b Higher is better.
Performance Analysis: The results shown above demonstrate competitive performance across all three metrics. Grid search consistently achieves the best solutions, benefiting from exhaustive evaluation of the discretized parameter space. However, the LLM-based optimization framework demonstrates far better optimization performance compared to conventional methods. Table 3 shows the number of iterations required for each optimization approach across different metrics.
Table 3: Optimization Performance Comparison
(a) Convergence Speed (Iterations to Converge)
Metric LLM IPOPT
Cost (\$/yr) 21 35
Yield (mol/yr) 26 81
Yield/Cost (mol/\$) 43 46
(b) Optimization Wall Time (Hours)
Metric LLM Grid Search
Cost (\$/yr) 0.17 10.50
Yield (mol/yr) 0.20 10.50
Yield/Cost (mol/\$) 0.33 10.50
The LLM framework consistently required fewer iterations than IPOPT. Most notably, for yield optimization, IPOPT required 81 iterations compared to only 26 for the LLM approach, which is a 3-fold reduction. The time efficiency gains over exhaustive methods are even more dramatic. Table 3 shows the LLM framework completing optimization in under 20 minutes across all metrics, compared to grid search’s fixed 10.5-hour requirement, which is a 31-fold speedup. While grid search must evaluate all 10,000 parameter combinations by design, the LLM framework converges in around 20-45 iterations through intelligent search strategy adaptation.
Despite requiring significantly fewer iterations and computational time, the LLM framework achieves competitive solution quality. IPOPT performs marginally better in cost minimization and yield-to-cost ratio optimization, while the LLM framework outperforms IPOPT in yield maximization. These modest performance differences demonstrate the framework’s feasibility for constraint-limited optimization problems, particularly given its substantial computational advantages. The framework’s ability to balance solution quality with computational efficiency makes it especially valuable for scenarios requiring rapid optimization or when computational resources are limited.
All computational benchmarks were performed on a MacBook Pro equipped with an Apple M1 chip (8-core, 16 GB RAM).
Reasoning-Guided Parameter Exploration: To examine the agent’s internal decision-making logic, we conducted a separate, additional trial with relaxed verbosity constraints. For cost minimization, the agent demonstrated sophisticated process understanding as shown below:
“Keep the reactor charge just inside the allowable window – the nearer-to-minimum H101 temperature ($\approx 830K$) cuts furnace duty and fuel. Push both flash drums to the warm end of their ranges; warmer condensers mean less refrigeration/brine duty. Relax the second-flash pressure drop toward the mechanical minimum (–30 kPa). Throttling is essentially "free," so choosing the smallest $\Delta P$ avoids excessive vapour generation that would otherwise have to be recompressed or condensed downstream.”
This reasoning demonstrates three key capabilities: correct identification of heating-cooling utility trade-offs, understanding of operating condition impacts on energy consumption, and application of domain-informed heuristics for parameter selection. The agent’s ability to articulate process engineering principles while making parameter adjustments distinguishes this approach from conventional numerical optimizers. While this local cost-minimization strategy may not directly lead to the global optimum, it provides efficient search directions that can be refined through subsequent iterations.
Table 4 demonstrates how the agent’s reasoning translates into systematic parameter adjustments during cost minimization. The optimization process begins with arbitrary user-provided initial guesses that may violate the autonomously generated constraints. As shown, the initial H101 temperature of 600 K violates the average generated range [827.2, 975.2] K, resulting in "Invalid" status until the ValidationAgent-SuggestionAgent loop produces the first feasible solution at iteration 1. The progression reveals the agent’s ability to systematically test different operating strategies and adapt based on performance feedback. Early parameter exploration shows varying cost performance as the agent tests different operating strategies, while subsequent iterations show the framework’s convergence toward more cost-effective operating conditions. By iteration 10, the agent identifies an optimal reactor temperature of 830 K that substantially reduces operating costs, with later iterations focusing on fine-tuning secondary variables while maintaining this temperature setting.
Table 4: Systematic parameter exploration during operating cost minimization.
Iteration H101 Temp (K) F101 Temp (K) F102 Temp (K) F102 $\Delta P$ (Pa) Cost (\$/yr)
User Initial Guess 600.0 325.0 375.0 -240000 Invalid
1 840.0 325.0 360.0 -180000 601715
2 830.0 330.0 363.0 -160000 617356
6 840.0 330.0 357.0 -175000 632627
10 830.0 325.0 360.0 -180000 579186
$\vdots$ $\vdots$ $\vdots$ $\vdots$ $\vdots$ $\vdots$
Methodological Advantages: The LLM approach does not require objective functions to be differentiable or continuous, potentially making it applicable to optimization problems with discontinuous or non-differentiable objectives. The scalability advantages become particularly evident when considering problem dimensionality. Grid search suffers from exponentially increasing computational demands as the number of decision variables grows, with our four-variable implementation already requiring over 10 hours wall time despite using only 10 discretization points per variable. While the LLM framework’s scalability with problem complexity remains to be validated, its current efficiency suggests potential advantages for complex optimization scenarios where intelligently guided parameter exploration is essential.
4 Conclusions
This work demonstrates that LLM agents can autonomously solve the constraint definition bottleneck that limits traditional optimization methods. By generating engineering-appropriate operating bounds from minimal process descriptions and iteratively optimizing within those bounds, our multi-agent framework enables process optimization in scenarios previously requiring extensive domain expertise or manual constraint specification.
The framework’s primary innovation lies in autonomous constraint generation combined with collaborative multi-agent optimization. By leveraging embedded domain knowledge to generate engineering-appropriate operating bounds for unit operations, the system enables optimization in scenarios where traditional methods cannot be directly applied. Combined with flexibility in handling non-differentiable objective functions, this capability enables optimization where traditional gradient-based methods may encounter difficulties. Validation on the HDA process demonstrated competitive performance against IPOPT and grid search across multiple objectives, with significant computational efficiency gains achieving convergence in under 20 minutes compared to over 10 hours for grid search. Notably, the framework’s reasoning capabilities distinguish it from conventional optimizers, with agents demonstrating sophisticated process understanding through explicit trade-off identification and domain-informed parameter selection strategies.
This work establishes a foundation for autonomous process optimization that could significantly reduce expertise barriers and enable wider application of advanced optimization techniques. The demonstrated feasibility on the HDA process provides a proof-of-concept for the multi-agent approach, opening pathways for extending functionality toward autonomous process modeling and flowsheet synthesis. Key opportunities for advancement include integration with process databases for better constraint generation, and development of approaches that combine LLM reasoning with conventional optimization methods for improved robustness.
5 Code Availability
All code used in this study is openly available at:
[…]
{suppinfo}
The Supporting Information includes multi-agent algorithm specifications, complete agent prompt templates, ContextAgent process understanding demonstrations, constraint generation approach comparisons, and detailed HDA process implementation including feed specifications, economic objective function formulations with utility cost coefficients, operational constraints, and IDAES modeling details.
6 Author Contributions
T.Z., S.B., J.O., C.L and A.B.F. conceived the project. T.Z. developed the core multi-agent optimization framework, and led the manuscript preparation. S.B. contributed to the development of the constraint generation modules, refined the core optimization algorithm, and assisted with manuscript writing. J.O. provided key conceptual input on the design of the overall framework and optimization strategy. C.L. contributed to the implementation and validation of the IDAES-based simulations. All authors reviewed and approved the final manuscript. A.B.F. supervised the project.